# Trust & Safety of LLMs and LLMs in Trust & Safety

### v1

@@ 2 Trust and Safety of LLMs
Large language models (LLMs) have emerged as a transformative force in artificial intelligence, demonstrating remarkable capabilities in natural language processing, text generation, and knowledge retrieval. However, the rapid advancement and widespread adoption of LLMs have also raised critical concerns regarding their trustworthiness and safety. These concerns stem from the potential for LLMs to generate biased or discriminatory outputs, spread misinformation, and be exploited for malicious purposes. Bender et al., 2021[BGMMS21]. highlight the inherent risks of stochastic parrots, emphasizing the potential for LLMs to perpetuate and amplify societal biases present in their training data. Similarly, Weidinger et al., 2021[WMR+21] delve into the ethical challenges posed by LLMs, including issues of fairness, accountability, and transparency.
@@ 2.1 Trust in LLMs
Furthermore, Bender et al. (2021) in ”On the Dangers of Stochastic Parrots: Can Language Models Be Too Big?” underscored the potential for LLMs to amplify societal biases present in their training data, calling for greater attention to fairness and inclusivity in LLM development to mitigate the risk of discriminatory outputs. Bender et al., 2021[BGMMS21] This highlights the ethical imperative to address bias and promote fairness in LLMs to prevent the perpetuation and amplification of harmful stereotypes. Similarly, Nadeem et al. (2021) in ”Stereotypes in Large Language Models” systematically investigated the presence of stereotypes related to gender, race, and religion in LLMs, providing empirical evidence of biased associations and highlighting the need for debiasing techniques. Nadeem et al., 2021[NBR21] This reinforces the need for ongoing research into debiasing techniques and the development of LLMs that are sensitive to social and cultural contexts.
@@ 2.2 Safety in LLMs
Ji et al. (2023) provide a comprehensive overview of LLM safety and trustworthiness challenges in their survey paper ”A Survey of Safety and Trustworthiness of Large Language Models.” This work examines various risks, including bias, toxicity, misinformation, and adversarial attacks, offering a valuable resource for researchers and practitioners. Ji et al., 2023[JXG+23]
Finally, ”Safety Prompts: a Systematic Review of Open Datasets for Evaluating and Improving Large Language Model Safety” by Scheurer et al. (2024) offers a crucial resource for researchers by systematically reviewing over 100 datasets specifically designed for LLM safety evaluation and improvement. This work helps navigate the growing landscape of safety datasets and identify areas for future development. Scheurer et al., 2024[SKP+24]
@@ 2.3 Key Findings
Below table provides a starting point for understanding how researchers are approaching the crucial task of measuring LLM trust and safety according to a survey paper from He et al. , 2023[SHW+24].
@@ 3.1 Health
Personalized Healthcare and Diagnostics: LLMs hold immense promise in analyzing patient data and delivering personalized recommendations, potentially improving access to healthcare and enabling early diagnosis. However, the dissemination of inaccurate information and potential privacy breaches are significant concerns. Liu et al., 2023[Liu23] provide a comprehensive analysis of this landscape, emphasizing the need for robust validation and safeguards to ensure accuracy and patient privacy.
Mental Health Support: LLMs offer a novel avenue for providing conversational support for individuals facing mental health challenges, increasing accessibility to resources, particularly for those who value anonymity or face barriers to traditional therapy. However, it is crucial to recognize that LLMs cannot replace human therapists, and the potential for misdiagnosis or inappropriate advice remains a concern. Vaidyam et al., 2023[Vai23] explore this evolving landscape, highlighting both the potential benefits and the need for careful implementation and oversight.
Drug Discovery and Development: LLMs are being explored for their potential to revolutionize drug discovery by analyzing vast amounts of biomedical data. This could accelerate the identification of promising drug candidates and lead to breakthroughs in disease treatment. However, inherent biases within the data could lead to inequitable outcomes, and the potential for misuse in developing harmful substances raises ethical concerns. Lo et al., 2023[Lo23] delve into the possibilities and challenges of employing LLMs in this critical field.
Trust and Safety: Ensuring the trust and safety of LLM outputs is paramount in healthcare applications. Choudhury et al., 2024[CS24] explore the potential impact of LLMs on user trust and the deskilling of healthcare professionals, emphasizing the need for cautious integration and user scrutiny of LLM outputs. Singh et al., 2024[SBM24] focus on the safety of medical LLMs, advocating for a shift towards prioritizing patient well-being and proactively mitigating potential risks. Szalai et al., 2023[Sza23] examine the trustworthiness of LLMs for health literacy efforts, stressing the importance of evaluating the reliability of LLM outputs and establishing trust in their application.
@@ 3.2 Finance
Fraud Detection and Security: LLMs offer the potential to bolster security measures and mitigate losses through advanced fraud detection by analyzing transaction patterns. However, the risk of false positives and the potential for malicious actors to exploit the system are critical concerns. Kumar et al., 2023[Kum23] highlight the dual-use nature of LLMs, demonstrating how they can be used to generate sophisticated financial fraud schemes.
@@
Automating Financial Reports: LLMs offer the potential to automate the generation of financial reports, reducing manual effort and improving efficiency. However, ensuring the accuracy and reliability of these reports is crucial to avoid significant financial and legal implications. Deng et al. (2023)[DWZ+23] explore the use of LLMs for generating financial reports, highlighting the need for careful validation and human oversight.
@@ 3.4 Best practices of using LLM in Trust and Safety field
      Furthermore, it is crucial to assess the dataset for redundancy and eliminate duplicate data points in the input to prevent inflated reasoning by Large Language Models (LLMs). Deduplication is especially important when the dataset is compiled from multiple sources and the same item appears in a redundant manner.You et al., 2024[YLF24], Lee et al., 2023[LSL+23], Abbas et al., 2023[AGLS23]
      Subsequently, the quality of the golden dataset needs to be evaluated in terms of annotation pollution. If the labeled data is not genuinely accurate, it can significantly influence the outcome of the LLM and diminish the classification resolution. Hence, to enhance the trustworthiness of LLMs, an additional layer of effort is required to measure the accuracy of the annotation data.Pavlick et a;., 2016[PCB16],Ribeiro et a;., 2018[RPCC18]
@@
      Therefore, analogous to the Chain of Thought (CoT) approach, a chain of LLMs should be employed to detect true positives and negatives, followed by an additional LLM at the end of the chain to identify harmful negative cases that should not be passed through the model.Jang et al., 2024[JLLK24],Lin et al., 2021[LHE21]
@@ 3.5 Emerging issues and solution
Jailbreak Attacks: LLMs are often trained with safety guidelines to prevent the generation of harmful or offensive content. Jailbreak attacks aim to circumvent these safety measures by exploiting loopholes or inconsistencies in the model’s training data or architecture.Xu et al., 2023[XLT+23] These attacks utilize cleverly crafted prompts to elicit undesirable outputs, such as:
@@
Red-teaming: Red-teaming, a security assessment technique simulating adversarial attacks, emerges as a crucial methodology for identifying and mitigating these risks. It necessitates a multifaceted strategy that encompasses a variety of techniques and methodologies aimed at uncovering vulnerabilities and evaluating the robustness of existing safety mechanisms. This process involves a deep dive into adversarial prompt engineering, wherein malicious prompts are meticulously crafted to exploit potential weaknesses, trigger unintended behaviors, or bypass the established safety filters. This encompasses various tactics, including prompt injection, where malicious code or commands are subtly embedded within user prompts to manipulate the LLM’s operational framework. Ganguli et al., 2022[GHLS22], Perez et al., 2022[PR22]. Furthermore, jailbreaking techniques are employed, utilizing carefully constructed prompts designed to circumvent safety guidelines and elicit harmful or biased outputs [2]. Additionally, data poisoning, the introduction of malicious data into the training dataset, represents another avenue for adversaries to influence the LLM’s responses and introduce vulnerabilities [1].

### v2

@@ 2 Trust and Safety of LLMs
Large language models (LLMs) have emerged as a transformative force in artificial intelligence, demonstrating remarkable capabilities in natural language processing, text generation, and knowledge retrieval. However, the rapid advancement and widespread adoption of LLMs have also raised critical concerns regarding their trustworthiness and safety. These concerns stem from the potential for LLMs to generate biased or discriminatory outputs, spread misinformation, and be exploited for malicious purposes. [BGMMS21]. highlight the inherent risks of stochastic parrots, emphasizing the potential for LLMs to perpetuate and amplify societal biases present in their training data. Similarly, [WMR+21] delve into the ethical challenges posed by LLMs, including issues of fairness, accountability, and transparency.
@@ 2.1 Trust in LLMs
Furthermore, Bender et al. (2021) in ”On the Dangers of Stochastic Parrots: Can Language Models Be Too Big?” underscored the potential for LLMs to amplify societal biases present in their training data, calling for greater attention to fairness and inclusivity in LLM development to mitigate the risk of discriminatory outputs. Bender et al., 2021[BGMMS21] This highlights the ethical imperative to address bias and promote fairness in LLMs to prevent the perpetuation and amplification of harmful stereotypes. Similarly, Nadeem et al. (2021) in ”Stereotypes in Large Language Models” systematically investigated the presence of stereotypes related to gender, race, and religion in LLMs, providing empirical evidence of biased associations and highlighting the need for debiasing techniques. Shrawgi et al., 2024[SRSD24] This reinforces the need for ongoing research into debiasing techniques and the development of LLMs that are sensitive to social and cultural contexts.
@@ 2.2 Safety in LLMs
Ji et al. (2023) provide a comprehensive overview of LLM safety and trustworthiness challenges in their survey paper ”A Survey of Safety and Trustworthiness of Large Language Models.” This work examines various risks, including bias, toxicity, misinformation, and adversarial attacks, offering a valuable resource for researchers and practitioners. Huang et al., 2023[HRH+23]
Finally, ”Safety Prompts: a Systematic Review of Open Datasets for Evaluating and Improving Large Language Model Safety” by Scheurer et al. (2024) offers a crucial resource for researchers by systematically reviewing over 100 datasets specifically designed for LLM safety evaluation and improvement. This work helps navigate the growing landscape of safety datasets and identify areas for future development. Rottger et al., 2025[RPVH25]
@@ 2.3 Key Findings
Below table provides a starting point for understanding how researchers are approaching the crucial task of measuring LLM trust and safety according to a survey paper from Huang et al. , 2024[HSW+24].
@@ 3.1 Health
Personalized Healthcare and Diagnostics: LLMs hold immense promise in analyzing patient data and delivering personalized recommendations, potentially improving access to healthcare and enabling early diagnosis. However, the dissemination of inaccurate information and potential privacy breaches are significant concerns. Zabir et al., 2023[NP24] provide a comprehensive analysis of this landscape, emphasizing the need for robust validation and safeguards to ensure accuracy and patient privacy.
Mental Health Support: LLMs offer a novel avenue for providing conversational support for individuals facing mental health challenges, increasing accessibility to resources, particularly for those who value anonymity or face barriers to traditional therapy. However, it is crucial to recognize that LLMs cannot replace human therapists, and the potential for misdiagnosis or inappropriate advice remains a concern. Vaidyam et al., 2019[VWH+19] explore this evolving landscape, highlighting both the potential benefits and the need for careful implementation and oversight.
Drug Discovery and Development: LLMs are being explored for their potential to revolutionize drug discovery by analyzing vast amounts of biomedical data. This could accelerate the identification of promising drug candidates and lead to breakthroughs in disease treatment. However, inherent biases within the data could lead to inequitable outcomes, and the potential for misuse in developing harmful substances raises ethical concerns. Zheng et al., 2024[ZKY+24] delve into the possibilities and challenges of employing LLMs in this critical field.
Trust and Safety: Ensuring the trust and safety of LLM outputs is paramount in healthcare applications. Choudhury et al., 2024[CS24] explore the potential impact of LLMs on user trust and the deskilling of healthcare professionals, emphasizing the need for cautious integration and user scrutiny of LLM outputs. Han et al., 2024[HKAL24] focus on the safety of medical LLMs, advocating for a shift towards prioritizing patient well-being and proactively mitigating potential risks. Szalai et al., 2023[Sza23] examine the trustworthiness of LLMs for health literacy efforts, stressing the importance of evaluating the reliability of LLM outputs and establishing trust in their application.
@@ 3.2 Finance
Fraud Detection and Security: LLMs offer the potential to bolster security measures and mitigate losses through advanced fraud detection by analyzing transaction patterns. However, the risk of false positives and the potential for malicious actors to exploit the system are critical concerns. Kumar et al., 2023[Pil23] highlight the dual-use nature of LLMs, demonstrating how they can be used to generate sophisticated financial fraud schemes.
@@
Automating Financial Reports: LLMs offer the potential to automate the generation of financial reports, reducing manual effort and improving efficiency. However, ensuring the accuracy and reliability of these reports is crucial to avoid significant financial and legal implications.
@@ 3.4 Best practices of using LLM in Trust and Safety field
      Furthermore, it is crucial to assess the dataset for redundancy and eliminate duplicate data points in the input to prevent inflated reasoning by Large Language Models (LLMs). Deduplication is especially important when the dataset is compiled from multiple sources and the same item appears in a redundant manner.You et al., 2024[YF24], Tirumala et al., 2023[TSAM23], Abbas et al., 2023[AGLS23]
      Subsequently, the quality of the golden dataset needs to be evaluated in terms of annotation pollution. If the labeled data is not genuinely accurate, it can significantly influence the outcome of the LLM and diminish the classification resolution. Hence, to enhance the trustworthiness of LLMs, an additional layer of effort is required to measure the accuracy of the annotation data.Putra et a;., 2024[PSS24],Klie et a;., 2024[KHKN24]
@@
      Therefore, analogous to the Chain of Thought (CoT) approach, a chain of LLMs should be employed to detect true positives and negatives, followed by an additional LLM at the end of the chain to identify harmful negative cases that should not be passed through the model.Jang et al., 2024[SYY24],Lin et al., 2021[LHE21]
@@ 3.5 Emerging issues and solution
Jailbreak Attacks: LLMs are often trained with safety guidelines to prevent the generation of harmful or offensive content. Jailbreak attacks aim to circumvent these safety measures by exploiting loopholes or inconsistencies in the model’s training data or architecture.Liu et al., 2024[LDX+24] These attacks utilize cleverly crafted prompts to elicit undesirable outputs, such as:
@@
Red-teaming: Red-teaming, a security assessment technique simulating adversarial attacks, emerges as a crucial methodology for identifying and mitigating these risks. It necessitates a multifaceted strategy that encompasses a variety of techniques and methodologies aimed at uncovering vulnerabilities and evaluating the robustness of existing safety mechanisms. This process involves a deep dive into adversarial prompt engineering, wherein malicious prompts are meticulously crafted to exploit potential weaknesses, trigger unintended behaviors, or bypass the established safety filters. This encompasses various tactics, including prompt injection, where malicious code or commands are subtly embedded within user prompts to manipulate the LLM’s operational framework. [PHS+22], Perez et al., 2022[PR22]. Furthermore, jailbreaking techniques are employed, utilizing carefully constructed prompts designed to circumvent safety guidelines and elicit harmful or biased outputs [2]. Additionally, data poisoning, the introduction of malicious data into the training dataset, represents another avenue for adversaries to influence the LLM’s responses and introduce vulnerabilities [1].